OpenStudy (_excusemyfrench._):

A costume designer will buy between 5 and 12 hats for the actors in a play. Each hat costs $7. the function c(h)= 7h models the total amount if money the costume designer spends on h hats. What is the practical range of the function? A. all real numbers from 5 to 12, inclusive B. all real numbers from 35 to 84, inclusive C. all integers from 5 to 12 inclusive D. all multiples of 7 between 35 and 84 inclusive

OpenStudy (decarr432):

Are there answer choices given?

OpenStudy (_excusemyfrench._):

yes i forgot to put them in the question. 1 sec

OpenStudy (decarr432):

take your time

OpenStudy (_excusemyfrench._):

okay the answer choices are posted

OpenStudy (decarr432):

Okay what is the difference between real numbers and integers do you know?

OpenStudy (_excusemyfrench._):

um not really

OpenStudy (decarr432):

okay A whole number is just a number itself no dec. and no frac. A interger is an integer is a number itself that cannot be written as a dec. or frac. do you see what I see?

OpenStudy (decarr432):

They are just the same thing

OpenStudy (_excusemyfrench._):

well..that's kinda confusing lol

OpenStudy (decarr432):

Think of it as the same thing because thats what they are the exact same thing

OpenStudy (_excusemyfrench._):

Okay

OpenStudy (decarr432):

So what do you know now review

OpenStudy (_excusemyfrench._):

Real numbers and integers are the same thing

OpenStudy (decarr432):

Good now take c(h)= 7h and substitute numbers in start at five and go to twelve c(5)=7(5)=35

OpenStudy (_excusemyfrench._):

so would that be B.?

OpenStudy (decarr432):

I believe so but I am not 100% on it sorry

OpenStudy (anonymous):

C. The possible hat count is 5, 6, 7, ... 12, ie; there are an integral number of hats to be purchased.

OpenStudy (pawanyadav):

C(h)=7h And h=5,6,7,8.........,12 So range=5×7,6×7,...............,12×7 =35,42,.................,84 So all multiple of 7 from 35 to 84 inclusive

OpenStudy (pawanyadav):

So, what's the answer?

OpenStudy (_excusemyfrench._):

D. thank you!!

OpenStudy (pawanyadav):

You are welcome